Two insulating small spheres are rubbed against each other and placed 1cm apart.If they attract each other with a force of 0.1 N how many electrons weretransfered from 1 sphere to the otherduring rubbing

Grade:12

4 Answers

askiitian.expert- chandra sekhar
14 years ago

Hi Pallavi,

F= 0.1 N = 9 * 109 * q2/r2

q2= 0.1 * 0.012/(9 * 109)

q2= 10/9 * 10-15

no of electrons transferred = q/1.6 * 10-19= 3.3333*10-16/1.6 * 10-19

                                                             =2.08 *103

s maheshwara srinivas
9 years ago
2*1011 electrons
Anjali Singh
9 years ago
 

F= 0.1 N = 9 * 109 * q2/r2

 

q2= 0.1 * 0.012/(9 * 109)

 

q2= 10/9 * 10-15

 

q= 3.34* 10-8
 

no of electrons transferred = q/1.6 * 10-19= 3.34* 10-8/1.6 * 10-19

 

                                                             =2 *1011 (Approx.)

Krish Gupta
3 years ago
 

Hi Pallavi,

F= 0.1 N = 9 * 109 * q2/r2

q2= 0.1 * 0.012/(9 * 109)

q2= 10/9 * 10-15

 

F= 0.1 N = 9 * 109 * q2/r2

 

q2= 0.1 * 0.012/(9 * 109)

 

q2= 10/9 * 10-15

 

q= 3.34* 10-8
 

no of electrons transferred = q/1.6 * 10-193.34* 10-8/1.6 * 10-19

 

                                                             =2 *1011 (Approx.)

Helps !